A fisherman passes the time by counting waves passing under his boat. He notices that his boat rises and falls 8 times in 65 seconds. He counts 6 wave crests between his boat and a buoy located 54 m away. What is the speed of the water waves?


1
Expert's answer
2021-02-24T16:59:12-0500

"v=\frac s t,\\
\nu=\frac {n}{t_1},\\
t=\frac {k}{\nu},\\
v=\frac{sn}{kt_1},\\
v=\frac{54\cdot 8}{6\cdot 65}\approx 1.1~\frac{\text{m}}{\text{s}}."